Problem 8.10 Debra King wants to invest in four-year bonds that are currently priced at $898.49. These bonds have a coupon rate of 6.0 percent and make semiannual coupon payments. What is the current market yield on this bond? (Round intermediate calculations to 5 decimal places, e.g. 1.25145 and final answer to 2 decimal places, e.g. 15.25%.) 

Current market yield = _______%
Answer #1

Answer:

Given

Current Price P=$898.49

Face Value F=$1000

Coupon rate =6%

Semi annual coupon rate C=6%/2=3%

Number of Years =4 years

Number of Semi annuals N=4*2=8

Let r be the semi annual current market yield.

so

P=C*F*(1-(1+r)^-N)r + F/(1+r)^N

898.49=3%*1000*(1-(1+r)^-8)/r + 1000/(1+r)^8

Solving for r we get r=4.54%

Annual current market yield =2*4.54%=9.08%